True or False:

A half mole of CO2 and a half mole of CS2 have the same number of molecules.

and

the complete combustion of butane, C4H10 produces CO2 and H2O.
Answer #1

1 mole of CO2   = 6.023*10^23 molecules

0.5(half) mole of Co2 = 6.023*10^23*0.5/1 = 3.0115*10^23 molecules

1 mole of CS2 = 6.023*10^23 molecules

0.5(half) mole of CS2 = 6.023*10^23*0.5/1 = 3.0115*10^23 molecules

True >>>>answer

The complete combustion of butane, C4H10 produces CO2 and H2O.

2C4H10(g) + 13O2(g) -------------> 8CO2(g)   + 10H2O(g)

true >>>>answer